# UniADC: A Unified Framework for Anomaly Detection and Classification

### v1

@@ Abstract
In this paper, we introduce the task of unified anomaly detection and classification, which aims to simultaneously detect anomalous regions in images and identify their specific categories. Existing methods typically treat anomaly detection and classification as separate tasks, thereby neglecting their inherent correlation, limiting information sharing, and resulting in suboptimal performance. To address this, we propose UniADC, a Unified Anomaly Detection and Classification model that can effectively perform both tasks with only a few or even no anomaly images. Specifically, UniADC consists of two key components: a training-free controllable inpainting network and a multi-task discriminator. The inpainting network can synthesize anomaly images of specific categories by repainting normal regions guided by anomaly priors, and can also repaint few-shot anomaly samples to augment the available anomaly data. The multi-task discriminator is then trained on these synthesized samples, enabling precise anomaly detection and classification by aligning fine-grained image features with anomaly-category embeddings. We conduct extensive experiments on three anomaly detection and classification datasets, including MVTec-FS, MTD, and WFDD, and the results demonstrate that UniADC consistently outperforms existing methods in anomaly detection, localization, and classification. The code is available at https://github.com/cnulab/UniADC.
1 Introduction
@@
Image anomaly detection aims to train models to detect and localize anomalous regions within images. It has numerous practical applications and has attracted growing research interest in recent years [37, 40, 5, 22]. Anomaly classification further categorizes detected anomalies, which aids in assessing anomaly severity and facilitating root-cause analysis. Existing studies [23, 13] typically treat anomaly classification as a downstream task of anomaly detection, in which the detected anomalous regions are cropped into sub-images and then fed into a separate classification model. This two-stage pipeline ignores the inherent connection between anomaly detection and classification, and hinders information sharing across the two tasks. Moreover, missed and over detections during the detection stage, as well as the difficulty in selecting appropriate anomaly score thresholds, pose a series of challenges for subsequent classification. As a result, such methods often suffer from high implementation complexity and suboptimal performance, limiting their practicality in real-world scenarios.
@@
To address the above tasks, we propose UniADC, a unified model for anomaly detection and classification, which tackles the challenge of scarce anomalous data through controllable image inpainting. It consists of two key components: a training-free controllable inpainting network and a multi-task discriminator. The controllable inpainting network consists of a pre-trained latent diffusion model [26] and an inpainting control network [44], enabling controllable inpainting guided by either anomaly priors or anomaly samples. Specifically, anomaly prior-guided controllable inpainting synthesizes category-specific anomalous images by repainting normal regions according to the provided prior knowledge of anomalies, whereas anomaly sample-guided controllable inpainting refines and repaints few-shot anomalous samples to enhance data diversity. In addition, we propose a category consistency selection strategy to filter synthetic anomaly samples that are highly consistent with the target category. The multi-task discriminator is trained on these synthetic anomaly samples and leverages the powerful generalization capability of the existing vision-language models such as CLIP [25] or dino.txt [16, 30] to achieve end-to-end anomaly detection and classification. As shown in Fig. 2, UniADC comprehensively outperforms existing methods in anomaly detection, localization, and classification, particularly strengthening the reliability of defect categorization. Our main contributions are summarized as follows:
@@
  • We propose a controllable inpainting network that can generate category-specific anomaly samples conditioned on either anomaly prior or few-shot anomaly images, without additional training. This enables broad applicability across various anomaly detection and classification tasks, serving as an effective alternative to existing anomaly synthesis methods.
@@
Figure 3: Overview of the proposed UniADC pipeline, which consists of a training-free controllable inpainting network and a multi-task discriminator. The controllable inpainting network supports two modes: anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting, enabling the generation of category-specific anomaly samples under different settings. The multi-task discriminator aligns fine-grained image features with anomaly category embeddings for accurate anomaly detection and classification.
2 Related Work
Few-shot Anomaly Detection. Few-shot anomaly detection methods train models using only a small number of normal samples and often rely on the general knowledge of foundation models to overcome the challenges posed by limited data. WinCLIP [14] employs the CLIP model [25] and aligns image patches with normal and anomalous prompts via sliding windows, enabling anomaly detection and fine-grained anomaly localization. PromptAD [20] introduces prompt learning to enhance the alignment process and refine decision boundaries. AnomalyGPT [8] feeds the query images and their anomaly scores into an LLM for anomaly inspection, eliminating the dependency on decision thresholds. It is worth noting that some existing methods also incorporate certain prior knowledge of anomalies to improve anomaly detection. For example, PromptAD [20] employs textual descriptions of each anomaly category to improve detection accuracy. AnomalyGPT [8] predefines the expected appearance of query images and potential anomaly types. AnoGen [9] constrains the size range of synthesized anomalies according to their types. In contrast, the proposed UniADC leverages anomaly prior knowledge for both anomaly detection and classification, providing a functional extension to existing methods.
@@ 2 Related Work
Anomaly Classification. Anomaly classification aims to identify the categories of detected anomalous regions, which remains highly challenging due to the scarcity of anomaly samples. ZipAF [23] employs the AlphaCLIP model [35] to extract region-contextual anomaly features and adopts a zero-initialized projection to align query features with cached anomaly features for few-shot anomaly classification. However, it does not incorporate the anomaly detection process, which may result in degraded performance in practical scenarios. AnomalyNCD [13] integrates novel class discovery methods into the existing anomaly detection pipeline for anomaly detection and clustering. Recent work [28] proposes a zero-shot multi-type anomaly detection method, MultiADS, which trains the model to learn cross-domain anomaly priors for anomaly classification. However, due to the lack of targeted training, this method struggles to identify domain-specific anomaly types, such as identifying highly similar defects within a specific product [33, 12, 43].
@@
Anomaly Synthesis. Existing anomaly synthesis methods can generally be classified into zero-shot and few-shot approaches, depending on the availability of real anomaly data during the synthesis process [18]. Zero-shot approaches generate anomaly samples relying on predefined data augmentation rules [19, 29], noise injection [38, 46], or textual descriptions [34], without access to real anomaly samples. However, the synthesized anomalies may exhibit a distribution shift from real-world cases, limiting their effectiveness. In contrast, few-shot approaches [7, 10, 15, 32] aim to enhance data diversity by augmenting a limited number of real anomaly samples. To the best of our knowledge, UniADC is the first method capable of synthesizing anomalies under both zero-shot and few-shot settings, thereby significantly enhancing its applicability across a wide range of anomaly detection and classification tasks.
3 Method
@@
In this section, we present our proposed anomaly detection and classification model, UniADC, with its overall pipeline illustrated in Fig. 3. We first provide the problem definition, then describe the main functionalities and implementation details of the controllable inpainting network, including anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting. Finally, we introduce the multi-task discriminator of UniADC.
3.1 Problem Definition
@@
3.2 Anomaly Prior-Guided Controllable Inpainting
This module synthesizes category-specific anomaly samples based on provided anomaly prior, making it suitable for zero-shot anomaly detection and classification tasks. The anomaly prior includes the shape, size, and textual description of each anomaly category. Shape and size are used to generate the anomaly mask, while the textual description serves as a prompt to guide the synthesis of the desired anomalous appearance. Specifically, we define eight types of candidate masks to simulate various anomaly shapes, including Rectangle, Line, Polygon, Ellipse, Hollow Ellipse, Random Brush, Perlin Noise, and Foreground Mask. Except for Foreground Mask, each type is generated at three scales: Large, Medium, and Small. We use specific mask types for mask generation based on the given anomaly shape and size. For example, we specify the shape of the anomaly category “Hole” as Ellipse, which will generate elliptical masks of arbitrary sizes. When both shape and size priors are unavailable for an anomaly category, all mask types are used for mask generation.
For a given normal image $X$ and prior knowledge from anomaly category $y$, we first generate an anomaly mask $M^{y}$ under the shape-size constraints. Then, the latent diffusion model encodes the image $X$ and the text prompt $p^{y}$ into latent variables $z=\mathbb{E}_{v}(X)$ and $z^{p}=\mathbb{E}_{t}(p^{y})$, where $\mathbb{E}_{v}(\cdot)$ and $\mathbb{E}_{t}(\cdot)$ denote the vision and text encoders, respectively. We define the number of forward diffusion steps as $T^{\prime}=T\cdot\gamma$, where $T$ denotes the total number of diffusion steps in the original latent diffusion model, and $\gamma\in(0,1]$ controls the noise strength. A larger $\gamma$ injects more noise into the latent variable, resulting in synthesized anomalies that deviate further from the original image distribution and exhibit more visually prominent anomalous patterns. To ensure spatial consistency between synthesized anomalous and the anomaly mask, we use an inpainting control network to precisely control the locations of the synthesized anomalies. Specifically, at each time step $t\in\{1,2,\cdots,T^{\prime}\}$, the inpainting control network $\phi$ predicts the conditioning variable $z^{m}=\phi(z_{t},t,z,M^{y})$ of noisy latent $z_{t}$ for controlling the denoising process:
@@ 3.2 Anomaly Prior-Guided Controllable Inpainting
where $\Psi$ denotes the latent diffusion model, and the conditioning variables $z^{p}$ and $z^{m}$ respectively control the semantics and spatial positions of the synthesized anomalies. The inpainting control network ensures that the denoised latent representation $z_{0}$ remains consistent with the original latent $z$ outside the masked region, while repainting only the masked region to align with the prompt embedding $z^{p}$. Finally, the denoised latent representation $z_{0}$ is decoded by decoder $\mathbb{D}(\cdot)$ into the anomaly image $X^{y}=\mathbb{D}(z_{0})$. In our implementation, we adopt Stable Diffusion v1.5 [26] as the latent diffusion model and BrushNet [17] as the inpainting control network. Both models are trained on high-quality and diverse data, enabling UniADC to generalize effectively across diverse anomaly synthesis scenarios.
However, the above synthesis process does not ensure the semantic alignment between anomaly masks and their corresponding textual descriptions, which may lead to inconsistencies between the synthesized anomaly samples and their intended categories. For example, in the case of “a broken transistor lead”, the synthesized anomaly should be located in the lead region, rather than randomly elsewhere. To this end, we propose a category consistency selection strategy to further filter synthetic anomalous images with high category consistency, thereby eliminating noisy samples introduced by anomaly synthesis. Given an anomaly category $y$, we first generate a mini-batch of synthesized samples using diverse anomaly masks, denoted as $\mathcal{S}^{y}=\{(X_{b}^{y},M_{b}^{y})\}_{b=1}^{B}$, where $B$ is the mini-batch size. We then adopt the AlphaCLIP model [35] to evaluate the semantic consistency between the synthesized image-mask pairs and the textual description of the anomaly category. For the synthetic anomaly sample $(X^{y}_{b},M^{y}_{b})$, we calculate its category matching score as:
@@
where $\psi_{v}$ and $\psi_{t}$ denote the vision and text encoders of AlphaCLIP, respectively, and $\langle\cdot,\cdot\rangle$ denotes the inner product. The AlphaCLIP model aligns the image region features guided by the masks with the text embeddings. If the synthetic anomalous region fails to match the corresponding anomaly description, such as being placed in the wrong location, the matching score decreases. In the end, we select the synthetic anomaly sample with the highest matching score in the mini-batch $\mathcal{S}^{y}$ for discriminator training.
Table 1: Comparison of UniADC with alternative methods under zero-shot anomaly detection and classification settings. $\mathcal{y}$ denotes methods that use anomaly priors.
$K_{n}\;\;$ Method MVTec-FS [23] MTD [12] WFDD [3]
1 InCTRL [49] 92.36 - - - - 70.34 - - - - 96.66 - - - -
2 InCTRL [49] 93.01 - - - - 72.07 - - - - 97.36 - - - -
4 InCTRL [49] 93.62 - - - - 73.23 - - - - 97.38 - - - -
3.3 Anomaly Sample-Guided Controllable Inpainting
@@
Anomaly sample-guided controllable inpainting aims to enrich the diversity of anomalies by repainting few-shot anomaly samples, making it well-suited for few-shot anomaly detection and classification tasks. Given a normal image $X$ and an anomalous sample $(X^{y},M^{y})$, we first crop the anomalous region from $X^{y}$ based on the mask $M^{y}$, and paste it at a random location on the normal image $X$, yielding a preliminary synthetic anomaly image $\hat{X}^{y}$ and its corresponding mask $\hat{M}^{y}$. To increase the shape variation, we apply identical data augmentations, such as affine transformations, to both the anomalous region and its mask. However, at this stage, the synthesized anomalies still suffer from limited diversity and suboptimal visual coherence. To improve diversity and image quality, we use the latent diffusion model along with an inpainting control network to repaint the pasted regions in $\hat{X}^{y}$. The repainting process is guided by simple text prompts (e.g., the name of the anomaly category) and follows the same diffusion and denoising procedures as described in anomaly prior-guided inpainting. The degree of diversity and deviation from the original anomaly is controlled by the noise factor $\gamma$. To ensure semantic consistency, we perform category consistency selection by computing the Structural Similarity Index (SSIM) [39] between the original and repainted anomalous sub-images. These sub-images include both the anomalous region and a fixed amount of surrounding normal context. We compute the SSIM-based category matching score for each sample and retain the highest-scoring synthetic anomaly sample for discriminator training.
Table 2: Comparison of UniADC with alternative methods under few-shot anomaly detection and classification settings.
$K_{n}$ $K_{a}$ Method MVTec-FS [23] MTD [12] WFDD [3]
3.4 Multi-task Discriminator
We train the multi-task discriminator on synthetic anomalous samples so that it learns to align fine-grained image features with anomaly category embeddings, enabling unified anomaly detection and classification. First, we represent each synthetic anomalous sample as a 4-tuple $(X,y,M_{d},M_{c})$, where $X$ is the synthesized anomalous image, $y$ is the anomaly category label, and $M_{d}$ and $M_{c}$ respectively denote the anomaly detection mask and classification mask, for a pixel location $(i,j)$ satisfying:

### v2

@@ Abstract
In this paper, we introduce a novel task termed unified anomaly detection and classification, which aims to simultaneously detect anomalous regions in images and identify their specific categories. Existing methods typically treat anomaly detection and classification as separate tasks, thereby neglecting their inherent correlations and limiting information sharing, which results in suboptimal performance. To address this, we propose UniADC, a model designed to effectively perform both tasks with only a few or even no anomaly images. Specifically, UniADC consists of two key components: a training-free Controllable Inpainting Network and an Implicit-Normal Discriminator. The inpainting network can synthesize anomaly images of specific categories by repainting normal regions guided by anomaly priors, and can also repaint few-shot anomaly samples to augment the available anomaly data. The implicit-normal discriminator addresses the severe challenge of the imbalance between normal and anomalous pixel distributions by implicitly modeling the normal state, achieving precise anomaly detection and classification by aligning fine-grained image features with anomaly-category embeddings. We conduct extensive experiments on three anomaly detection and classification datasets, including MVTec-FS, MTD, and WFDD, and the results demonstrate that UniADC consistently outperforms existing methods in anomaly detection, localization, and classification. The code is available at https://github.com/cnulab/UniADC.
I Introduction
Image anomaly detection aims to train models to detect and localize anomalous regions within images. It has numerous practical applications and has attracted growing research interest in recent years [1, 2, 3, 4]. Anomaly classification further categorizes detected anomalies, which aids in assessing anomaly severity and facilitating root-cause analysis. For instance, severe anomalies like scratches and breakage significantly compromise product quality, necessitating immediate production halts for inspection. Conversely, minor anomalies like stains or foreign matter have a negligible impact and are typically not classified as genuine defects. Existing studies [5, 6] typically treat anomaly classification as a downstream task of anomaly detection, in which the detected anomalous regions are cropped into sub-images and then fed into a separate classification model. This two-stage pipeline ignores the inherent connection between anomaly detection and classification, and hinders information sharing across the two tasks. Moreover, missed and over detections during the detection stage, as well as the difficulty in selecting appropriate anomaly score thresholds, pose a series of challenges for subsequent classification. As a result, such methods often suffer from high implementation complexity and suboptimal performance, limiting their practicality in real-world scenarios.
@@ I Introduction
To address the above tasks, we propose UniADC, a unified model for anomaly detection and classification, which tackles the challenge of scarce anomalous data through controllable image inpainting. It consists of two key components: a training-free controllable inpainting network and an implicit-normal discriminator. The controllable inpainting network consists of a pre-trained latent diffusion model [13] and an inpainting control network [14], enabling controllable inpainting guided by either anomaly priors or anomaly samples. Specifically, anomaly prior-guided controllable inpainting synthesizes category-specific anomalous images by repainting normal regions according to the provided prior knowledge of anomalies, whereas anomaly sample-guided controllable inpainting refines and repaints few-shot anomalous samples to enhance data diversity. In addition, we propose a category consistency selection strategy to filter synthetic anomaly samples that are highly consistent with the target category. The implicit-normal discriminator is trained on these synthetic anomaly samples and mitigates performance degradation caused by pixel-level distribution imbalance by disentangling the normal state from various anomaly concepts, thereby enabling robust anomaly detection and classification. Through the aforementioned innovative designs, UniADC comprehensively outperforms existing methods in anomaly detection, localization, and classification, as shown in Fig. 2, particularly strengthening the reliability of defect categorization. Our main contributions are summarized as follows:
@@
  • We propose a training-free controllable inpainting network that can generate category-specific anomaly samples conditioned on either anomaly prior or few-shot anomaly images. This enables broad applicability across various anomaly detection and classification tasks, serving as an effective alternative to existing anomaly synthesis methods.
  • We propose an implicit-normal discriminator that mitigates performance degradation arising from the long-tail effect of anomaly data by implicitly modeling normal concepts, enabling it to effectively align fine-grained image features with anomaly-category embeddings for end-to-end anomaly detection and classification.
@@
II Related Work
II-A Anomaly Detection
Image anomaly detection has gained widespread attention in recent years, leading to extensive research. Due to the scarcity of anomaly samples in real-world scenarios, most methods [17, 18, 19, 20, 21] follow an unsupervised paradigm, training solely on normal data and identifying anomalies that deviate from normal patterns during inference. However, in practical production scenarios, anomaly information is not entirely out of reach. Consequently, recent studies have explored integrating anomaly priors or limited anomaly samples into the detection pipeline to improve reliability. For example, PromptAD [8] employs textual descriptions of each anomaly category to improve detection accuracy. AnomalyGPT [9] predefines the expected appearance of query images and potential anomaly types to enhance the reasoning capability of the anomaly question answering model. AnoGen [22] integrates size priors of different anomaly types into the anomaly synthesis pipeline to ensure class consistency of the synthesized anomaly samples. BGAD [23] and NAGL [24] leverage a small number of anomaly samples to refine the model’s decision boundary, thereby bolstering its discriminative capability for more accurate detection. Motivated by these advances and real-world requirements, we systematically integrate diverse anomaly priors and limited anomaly samples into a unified anomaly detection and classification pipeline, and propose UniADC, a general framework that functionally extends existing methods.
II-B Anomaly Classification
@@
Anomaly classification aims to identify the categories of detected anomalous regions, which remains highly challenging due to the scarcity of anomaly samples. Despite its broad range of applications, this field has not yet been extensively explored. ZipAF [5] employs the AlphaCLIP model [25] to extract region-contextual anomaly features and adopts a zero-initialized projection to align query features with cached anomaly features for few-shot anomaly classification. However, it does not incorporate the anomaly detection process, which may result in degraded performance in practical scenarios. AnomalyNCD [6] integrates novel class discovery methods into the existing anomaly detection pipeline for anomaly detection and clustering. Recent work [26] proposes a zero-shot multi-type anomaly detection method, MultiADS, which trains the model to learn cross-domain anomaly priors for anomaly classification. However, due to the lack of targeted training, this method struggles to identify domain-specific anomaly types, such as identifying highly similar defects within a specific product [27, 15, 28].
II-C Anomaly Synthesis
@@
Existing anomaly synthesis methods can generally be classified into zero-shot and few-shot approaches, depending on the availability of real anomaly data during the synthesis process [29]. Zero-shot approaches generate anomaly samples relying on predefined data augmentation rules [30, 31, 32], noise injection [33, 34], or textual descriptions [35, 36], without access to real anomaly samples. However, the synthesized anomalies may exhibit a distribution shift from real-world cases, limiting their effectiveness. In contrast, few-shot approaches [37, 38, 39, 40] aim to enhance data diversity by augmenting a limited number of real anomaly samples. Despite significant progress in anomaly synthesis techniques, existing research remains primarily focused on anomaly detection. The category controllability and accuracy of synthesized anomaly samples still require further validation to ensure their feasibility for more challenging anomaly classification tasks.
Figure 3: Overview of the proposed UniADC pipeline, which consists of a training-free controllable inpainting network and an implicit-normal discriminator. The controllable inpainting network supports two modes: anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting, enabling the generation of category-specific anomaly samples under different settings. The implicit-normal discriminator aligns fine-grained image features with anomaly category embeddings for accurate anomaly detection and classification.
III Method
@@
In this section, we present our proposed anomaly detection and classification model, UniADC, with its overall pipeline illustrated in Fig. 3. We first provide the problem definition, then describe the main functionalities and implementation details of the controllable inpainting network, including anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting. Finally, we introduce the implicit-normal discriminator of UniADC.
III-A Problem Definition
@@
Figure 4: Examples of anomaly masks generated from the GAP-Lib.
III-B Anomaly Prior-Guided Controllable Inpainting
This module synthesizes category-specific anomaly samples based on provided anomaly priors, making it suitable for zero-shot anomaly detection and classification tasks. The anomaly prior includes the shape, size, and textual description of each anomaly category. Shape and size are used to generate the anomaly mask, while the textual description serves as a prompt to guide the synthesis of the desired anomalous appearance. Specifically, to map various anomaly shapes and sizes into explicit spatial constraints, we constructed a Geometric Anomaly Prototype Library (GAP-Lib). It comprises eight types of candidate masks, including Rectangle, Line, Polygon, Ellipse, Hollow Ellipse, Random Brush, Perlin Noise, and Foreground Mask. The first seven types are used to simulate various local anomalous regions, with each type defined in three sizes: Large, Medium, and Small, as illustrated in Fig. 4. Each mask size corresponds to a specific size range. Foreground Mask uses the object’s foreground as a mask to synthesize global anomalies. We use specific mask types for mask generation based on the given anomaly shape and size. For example, we specify the shape of the anomaly category “Hole” as Ellipse, which will generate elliptical masks of arbitrary sizes. When both shape and size priors are unavailable for an anomaly category, all mask types are used for mask generation.
For a given normal image $X$ and prior knowledge from anomaly category $y$, we first generate an anomaly mask $M^{y}$ utilizing the proposed GAP-Lib. Then, the latent diffusion model encodes the image $X$ and the text prompt $p^{y}$ into latent variables $z=\mathbb{E}_{v}(X)$ and $z^{p}=\mathbb{E}_{t}(p^{y})$, where $\mathbb{E}_{v}(\cdot)$ and $\mathbb{E}_{t}(\cdot)$ denote the vision and text encoders, respectively. We define the number of forward diffusion steps as $T^{\prime}=T\cdot\gamma$, where $T$ denotes the total number of diffusion steps in the original latent diffusion model, and $\gamma\in(0,1]$ controls the noise strength. A larger $\gamma$ injects more noise into the latent variable, resulting in synthesized anomalies that deviate further from the original image distribution and exhibit more visually prominent anomalous patterns. To ensure spatial consistency between synthesized anomalies and the anomaly mask, we use an inpainting control network to precisely control the locations of the synthesized anomalies. Specifically, at each time step $t\in\{1,2,\cdots,T^{\prime}\}$, the inpainting control network $\phi$ predicts the conditioning variable $z^{m}=\phi(z_{t},t,z,M^{y})$ of noisy latent $z_{t}$ for controlling the denoising process:
@@ III-B Anomaly Prior-Guided Controllable Inpainting
where $\Psi$ denotes the latent diffusion model, and the conditioning variables $z^{p}$ and $z^{m}$ respectively control the semantics and spatial positions of the synthesized anomalies. The inpainting control network ensures that the denoised latent representation $z_{0}$ remains consistent with the original latent $z$ outside the masked region, while repainting only the masked region to align with the prompt embedding $z^{p}$. Finally, the denoised latent representation $z_{0}$ is decoded by decoder $\mathbb{D}(\cdot)$ into the anomaly image $X^{y}=\mathbb{D}(z_{0})$. In our implementation, we adopt Stable Diffusion v1.5 [13] as the latent diffusion model and integrate BrushNet [41] as a plug-and-play control network for inpainting. This configuration enables UniADC to achieve highly accurate control over anomaly placement and image-mask consistency, bypassing the need for costly post-processing or mask refinement steps common in other anomaly synthesis methods [39, 35, 36].
However, the above synthesis process does not ensure the semantic alignment between anomaly masks and their corresponding textual descriptions, which may lead to inconsistencies between the synthesized anomaly samples and their intended categories. For example, in the case of “a broken transistor lead”, the synthesized anomaly should be located in the lead region, rather than randomly elsewhere. To this end, we propose a Category Consistency Selection (CCS) strategy to further filter synthetic anomalous images with high category consistency, thereby eliminating noisy samples introduced by anomaly synthesis. Given an anomaly category $y$, we first generate a mini-batch of synthesized samples using diverse anomaly masks, denoted as $\mathcal{S}^{y}=\{(X_{b}^{y},M_{b}^{y})\}_{b=1}^{B}$, where $B$ is the mini-batch size. We then adopt the AlphaCLIP model [25] to evaluate the semantic consistency between the synthesized image-mask pairs and the textual description of the anomaly category. For the synthetic anomaly sample $(X^{y}_{b},M^{y}_{b})$, we calculate its category matching score as:
@@
where $\psi_{v}$ and $\psi_{t}$ denote the vision and text encoders of AlphaCLIP, respectively, and $\langle\cdot,\cdot\rangle$ denotes the inner product. The AlphaCLIP model aligns the image region features guided by the masks with the text embeddings. If the synthetic anomalous region fails to match the corresponding anomaly description, such as being placed in the wrong location, the matching score decreases. In the end, we select the synthetic anomaly sample with the highest matching score in the mini-batch $\mathcal{S}^{y}$ for subsequent discriminator training.
III-C Anomaly Sample-Guided Controllable Inpainting
@@
Anomaly sample-guided controllable inpainting aims to enrich the diversity of anomalies by repainting few-shot anomaly samples, making it well-suited for few-shot anomaly detection and classification tasks. Given a normal image $X$ and an anomalous sample $(X^{y},M^{y})$, we first crop the anomalous region from $X^{y}$ based on the mask $M^{y}$, and paste it at a random location on the normal image $X$, yielding a preliminary synthetic anomaly image $\hat{X}^{y}$ and its corresponding mask $\hat{M}^{y}$. To increase the shape variation, we apply identical data augmentations, such as affine transformations, to both the anomalous region and its mask. However, at this stage, the synthesized anomalies still suffer from limited diversity and suboptimal visual coherence. To improve diversity and image quality, we use the latent diffusion model along with an inpainting control network to repaint the pasted regions in $\hat{X}^{y}$. The repainting process is guided by simple text prompts (e.g., the name of the anomaly category) and follows the same diffusion and denoising procedures as described in anomaly prior-guided inpainting. The degree of diversity and deviation from the original anomaly is controlled by the noise factor $\gamma$. To ensure semantic consistency, we perform category consistency selection by computing the Structural Similarity Index (SSIM) [42] between the original and repainted anomalous sub-images. These sub-images include both the anomalous region and a fixed amount of surrounding normal context. We compute the SSIM-based category matching score for each sample and retain the highest-scoring synthetic anomaly sample for discriminator training. Table I compares UniADC with other anomaly synthesis methods. To the best of our knowledge, UniADC is the first method capable of synthesizing anomalies under both zero-shot and few-shot settings. Moreover, it supports both diverse mask generation and fine-grained control over anomaly categories, thereby significantly enhancing its applicability across a wide range of anomaly detection and classification tasks.
TABLE I: Comparison between UniADC and alternative anomaly synthesis methods. Note: “training-free” specifically characterizes the anomaly synthesis stage, independent of the subsequent training of the detection model.
Method Training-free Zero-shot Generation Few-shot Generation Diversity Mask Controllable Anomaly Types
DRÆM [43]
RealNet [34]
AnoGen [22]
AnomalyDiff [38]
DualAnoDiff [39]
AnomalyAny [35]
UniADC (ours)
III-D Implicit-Normal Discriminator
Beyond the challenge of scarce anomalous samples, another core difficulty in anomaly detection and classification lies in the severe distribution imbalance between normal pixels and various categories of anomalous pixels. Conventional discriminative anomaly detection models [44, 8, 26] typically align image patch features with class embeddings corresponding to “normal” and “anomaly” concepts through adapter fine-tuning or learnable text prompts. However, for multi-class anomaly classification, the long-tail effect of the data is significantly exacerbated. This causes such methods to inevitably overfit to the massive volume of normal pixels, ultimately leading to performance degradation in both anomaly detection and classification. To mitigate this, we propose the Implicit-Normal Discriminator (IND). Our key insight is to eliminate the explicit category embedding for the normal state. Instead of treating “normal” as a distinct class, we identify normal regions by the absence of all potential anomaly types. This exclusion-based logic forces the model to focus exclusively on the manifold of diverse anomaly concepts, thereby substantially alleviating the bias induced by imbalanced data distributions. Specifically, the proposed implicit-normal discriminator is trained using synthetic anomalous samples to align fine-grained image features with anomaly category embeddings, thereby enabling unified anomaly detection and classification. First, we represent each synthetic anomalous sample as a 4-tuple $(X,y,M_{d},M_{c})$, where $X$ is the synthesized anomalous image, $y$ is the anomaly category label, and $M_{d}$ and $M_{c}$ respectively denote the anomaly detection mask and classification mask, for a pixel location $(i,j)$ satisfying: